### v1

# Training Verification-Friendly Neural Networks via Neuron Behavior Consistency

Abstract
Formal verification provides critical security assurances for neural networks, yet its practical application suffers from the long verification time. This work introduces a novel method for training verification-friendly neural networks, which are robust, easy to verify, and relatively accurate. Our method integrates neuron behavior consistency into the training process, making neuron activation states consistent across different inputs in a local neighborhood, reducing the number of unstable neurons and tightening the bounds of neurons thereby enhancing neural network verifiability. We evaluated our method using the MNIST, Fashion-MNIST, and CIFAR-10 datasets across various network architectures. The results of the experiment demonstrate that networks trained using our method are verification-friendly across different radii and different model architectures, whereas other tools fail to maintain verifiability as the radius increases. We also show that our method can be combined with existing methods to further improve the verifiability of networks.
Introduction
Neural networks are increasingly applied in safety-critical domains such as autonomous driving (Urmson and Whittaker 2008) and flight control (Julian, Kochenderfer, and Owen 2019). However, they frequently struggle with a lack of robustness, as even minor perturbations to their inputs can lead to incorrect predictions (Bu et al. 2022; Chen et al. 2021; Song et al. 2021; Chen et al. 2023, 2022b; Zhao et al. 2021; Chen et al. 2022a). This is unacceptable in safety-critical applications, where consistent performance is imperative. Thus, it is desirable to develop methods to systematically advance the robustness verification of neural networks.
Existing methods for analyzing robustness can be divided into two categories: empirical analysis through adversarial attacks and mathematical proof via formal verification. While adversarial attacks generate misleading examples, they merely demonstrate the presence of adversarial samples without affirming their absence. In contrast, formal verification ensures the correctness of neural networks using logical and mathematical methods, allowing us to formally verify the robustness of neural networks. This rigorous verification is essential for safety-critical systems.
Advanced formal verification tools (Wang et al. 2021; Zhang et al. 2022a; Bak 2021; Katz et al. 2017), typically employ branch-and-bound algorithms for neural network verification. At the beginning of verification, abstract interpretation (Gehr et al. 2018; Mirman, Gehr, and Vechev 2018; Zhang et al. 2021, 2023; Guo et al. 2021) is usually used to abstract neurons. If the properties of the network remain undetermined after using symbolic propagation (Singh et al. 2019) to calculate neuron boundaries and using MILP (Tjeng, Xiao, and Tedrake 2019; Tran et al. 2020; Zhang et al. 2022b; Zhang, Song, and Sun 2023; Zhang et al. 2024) or SMT methods (Ehlers 2017; Huang et al. 2017; Katz et al. 2017; Zhao et al. 2022; Liu et al. 2024) for constraint solving, further branching is required. The branching process involves enumerating the activation states of unstable neurons, whose activation status cannot be determined through bound calculations, to introduce additional constraints, thereby refining the abstraction. Typically, neural networks contain numerous unstable neurons, and exploring these combinations of activation states requires exponential time, which limits the widespread application of formal verification techniques in practice.
In addition to developing ever-more sophisticated methods for post-training verification, researchers have investigated the idea of training neural networks that are easier to verify, as known as verification-friendly neural networks. Ideally, a training method for verification-friendly neural networks must satisfy the following requirements. First, (accuracy) the resultant neural network must have an accuracy that is comparable to neural networks trained conventionally. Second, (robustness) the resultant neural network must have improved robustness, which could be measured using existing adversarial attacks. Third, (verifiability) it must be easier to verify using existing or dedicated neural network verification techniques, which can be measured using the effectiveness of selected neural network verification methods.
There are mainly two existing approaches to promoting the verification-friendliness of neural networks. One approach involves post-processing the network through methods that modify the weights of the networks (Xiao et al. 2019; Baninajjar, Rezine, and Aminifar 2023). Another involves altering the neural network design and training process with considerations for verification (Xiao et al. 2019; Narodytska et al. 2019; Xu et al. 2024). However, these methods still have limitations. The effectiveness of post-training methods is limited by the network itself. Moreover, certain post-training methods such as using MILP to make the network more sparse (Baninajjar, Rezine, and Aminifar 2023), are computationally expensive and may not be adaptable for large networks. The ReLU Stable methods, which introduce ReLU Stable (RS) loss (Xiao et al. 2019) to reduce the number of unstable neurons, depend on the bounds of neurons, and when the perturbation radius changes, it often causes these bounds to shift, affecting the verification efficiency of the network. Certified training, mainly based on the heavy Interval Bound Propagation (IBP) method (Mirman, Gehr, and Vechev 2018; Gowal et al. 2018; Zhang et al. 2019a; Xu et al. 2020) suffers from a long training time and the gradient explosion or vanish problem. Adversarial training (Madry et al. 2018; Zhang et al. 2019b; Ganin et al. 2016; Zhu et al. 2017) typically increase network complexity but do not contribute to improving network verifiability.
In this work, we introduce a straightforward yet effective training method that enhances the verifiability of neural networks by enforcing the consistency of neuronal behavior, which we refer to as neuron behavior consistency (NBC), throughout the training process as a regularization term. A neuron is called behavior consistent if its activation state remains the same in a given input neighborhood. By maximizing the consistency of neurons, the unstable neurons are decreased, reducing the search space of the verification process. NBC can also help to tight the the bounds of neurons, as fewer unstable neurons introduce less error during bound calculation algorithms. Our approach can be scaled to larger networks compared to MILP-based methods. Moreover, the core of our method lies in the consistency of neuronal behavior without relying on heavy IBP methods, making a reduction in training epochs and ensuring the trained network maintains verifiability across different perturbation radii.
We evaluate our method using Fashion-MNIST, MNIST, and CIFAR-10 datasets across different architectures at various perturbation radii. Our method oputperforms in stable neuron ratio and achieved up to a speedup of 450% in verification time. Importantly, our method accelerates the verification while preserving the accuracy of the model , which is not commonly achieved by existing methods. In summary, our contributions are as follows:
  • We introduce a method of training verification-friendly networks by integrating neuron behavior consistency.
  • We evaluated our methodology on three well-known datasets. Experiments show that networks trained using our method can maintain verification-friendly properties across different radii and different model architectures.
  • We demonstrate that our method can be combined with existing methods to further improve the verifiability of networks, especially in the case of large networks.
  • We show that our method accelerates the verification process while preserving model accuracy and robustness.
Preliminary
In this section, we introduce the background of neural network verification problems and the general branch and bound verification framework.
Neural Networks Verification Problems
Given a neural network $f:\mathbb{R}^{m_{\mathrm{in}}}\rightarrow\mathbb{R}^{m_{\mathrm{out}}}$, with $m_{\mathrm{in}}$ input neurons and $m_{\mathrm{out}}$ output neurons, the goal of the neural network verification problem is to determine whether the output of the network satisfies a set of output constraints $\mathcal{P}$ for all inputs that meet the input constraints $\mathcal{C}$, formally defined as:
Definition 1 (Neural Network Verification Problem).
The neural network verification problem $\langle f,\mathcal{C},\mathcal{P}\rangle$ is to determine whether:
$\forall\boldsymbol{x}\in\mathbb{R}^{m_{\mathrm{in}}},\boldsymbol{x}\in\mathcal{C}\Rightarrow f(\boldsymbol{x})\in\mathcal{P},$ (1)
where $\mathcal{C}\subseteq\mathbb{R}^{m_{\mathrm{in}}}$ is the input constraints and $\mathcal{P}\subseteq\mathbb{R}^{m_{\mathrm{out}}}$ is the output constraints.
We focus on the local robustness verification problem $\langle f,\mathcal{C}_{\epsilon}(\boldsymbol{x}),\mathcal{P}_{c}\rangle$, which checks if the classification result $c$ is robust to input perturbations $\varepsilon$ in the norm $l_{\infty}$, where the input constraints $\mathcal{C}_{\epsilon}(\boldsymbol{x})$ defined as $\{\boldsymbol{x}\in\mathbb{R}^{m_{\mathrm{in}}}\mid|\boldsymbol{x}-\boldsymbol{x}^{0}|_{\infty}\leq\varepsilon\}$ and the output constraints $\mathcal{P}_{c}$ defined as $\{\boldsymbol{y}\in\mathbb{R}^{m_{\mathrm{out}}}\mid\bigwedge_{i\neq c}\boldsymbol{y}_{i}-\boldsymbol{y}_{c}\leq 0\}$.
General Verification Framework
State-of-the-art methods for solving neural network verification problems (Wang et al. 2021; Zhang et al. 2022a; Katz et al. 2019; Bak 2021) are typically based on branch-and-bound algorithms, consisting of three critical components: constraint solving, bound calculation, and branch selection.
As shown in Figure 1, the verification process starts with the bound calculation, where the upper and lower bounds of neuron output are estimated under specified input constraints. If these bounds are sufficiently precise, the properties of the network can be directly verified. Due to non-linear activation functions such as ReLU, computing these bounds can be complex, which is a problem often addressed through neuron-wise abstraction (Singh et al. 2019; Bak 2021). This abstraction uses linear bounds to approximate neuron outputs, thereby simplifying the bound calculations.
When the bound calculation does not suffice to verify the network’s properties, constraint solving is employed. This process involves determining if the constraints of the abstracted network can be satisfied, using Linear Programming (LP) or Mixed Integer Linear Programming (MILP) methods. Constraints typically include the input constraints $\mathcal{C}$, the negated output constraints $\neg\mathcal{P}$ and the constraints of the abstracted network itself. If these constraints are UNSAT (unsatisfiable), the property holds, which proves that the network can not be successfully attacked within $\mathcal{C}$; otherwise, the returned counterexample must be examined. If the counterexample is not a false positive caused by over-approximation, a violation of the network’s properties is identified. Otherwise, it suggests that the network is over-abstracted, necessitating branch selection to refine the abstraction.
Branch selection selects an unstable neuron (whose activation state cannot be determined through bound calculations) and splits it into active and inactive branches. Identifying the state of a neuron entails establishing new constraints that refine the abstraction. Branch selection strategies critically affect verification efficiency. While branching may require exploration of all unstable neurons in worst-case scenarios, minimizing these neurons reduces the theoretical branching upper bound.
Figure 1: The Branch and Bound (BaB) verification process
Training Verification-Friendly Networks via Neuron Behavior Consistency
In this section, we propose our method, neuron behavior consistency, to train verification-friendly neural networks.
Given a network $f$ with parameters $\theta$ and an underlying distribution $\mathcal{D}$, the traditional training process aims to optimize the parameters $\theta$ to minimize the expected loss:
$\min_{\theta}\underset{(\boldsymbol{x},{y})\sim\mathcal{D}}{\mathbb{E}}\left({loss}(\boldsymbol{x},\boldsymbol{y})\right),$ (2)
where $(\boldsymbol{x},{y})$ denotes the input and target label sampled from the distribution $\mathcal{D}$, and $\boldsymbol{y}$ is the one-hot vector representation of the target label $y$, that is, a vector with a single 1 at the index of the target label and 0 elsewhere. The most common loss function is the cross-entropy function for classification tasks, defined as $\text{CE}(f(\boldsymbol{x}),\boldsymbol{y})=-\sum_{i}\boldsymbol{y}_{i}\log f(\boldsymbol{x})_{i}$.
The ordinary training objective does not impose constraints on neurons, potentially resulting in a large number of unstable neurons. We thus propose an alternative training objective that aims to maximize the consistency of neurons. A neuron is consistent if its activation state remains the same in a given input neighborhood. Formally, given an input $\boldsymbol{x}$ and a neighboring input $\boldsymbol{x}^{\prime}$, the consistency of the $j$-th neuron of the $i$-th layer $n^{(i)}_{j}$ is defined as:
${\textsc{NBC}}(n^{(i)}_{j},\boldsymbol{x},\boldsymbol{x}^{\prime})=\begin{cases}1,&\text{if }\mathrm{sign}(f^{(i)}(\boldsymbol{x})_{j})=\mathrm{sign}(f^{(i)}(\boldsymbol{x}^{\prime})_{j}),\\ 0,&\text{otherwise},\end{cases}$ (3)
where $f^{(i)}(\boldsymbol{x})_{j}$ denotes the pre-activation value of the $j$-th neuron of the $i$-th layer when fed with input $\boldsymbol{x}$. Intuitively, for any input within a given neighborhood, if the activation states of individual neurons are highly consistent, the calculated boundaries (upper and lower bounds) are more likely to be tight. This coherence may reduce the occurrence of unstable neurons in the neural network.
To maximize (minimize the negative) this consistency, we incorporate a regularization term into the optimization objective, which can be represented as:
$\min_{\theta}\underset{(\boldsymbol{x},{y})\sim\mathcal{D},\boldsymbol{x}^{\prime}\in\mathcal{C}_{\varepsilon}(\boldsymbol{x})}{\mathbb{E}}[\text{CE}(f(\boldsymbol{x}),\boldsymbol{y})-\beta\sum_{n_{i}\in\mathcal{N}}\textsc{NBC}(n_{i},\boldsymbol{x},\boldsymbol{x}^{\prime})],$ (4)
where $\beta$ is a hyperparameter that controls the importance of the regularization term, and $\mathcal{N}$ denotes the set of neurons in the network.
Adapting concepts from adversarial training, the loss function can be reformulated to maximize the minimal (minimize the negative minimal) consistency of neural behavior across different inputs within the neighborhood domain:
$l^{\textsc{NBC}}(\boldsymbol{x},{y})=\text{CE}(f(\boldsymbol{x}),\boldsymbol{y})-\beta\underset{\boldsymbol{x}^{\prime}\in\mathcal{C}_{\varepsilon}(\boldsymbol{x})}{\min}\sum_{n_{i}\in\mathcal{N}}\textsc{NBC}(n_{i},\boldsymbol{x},\boldsymbol{x}^{\prime}).$ (5)
The consistency metric presented in Equation 3 is a discrete measure that cannot be directly integrated into the loss function. Therefore, we employ a continuous metric to approximate the consistency of neural behavior across different inputs within the neighborhood domain, which is shown in Algorithm 1. This algorithm calculates the NBC between the neural network $f$ when fed with $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$.
Algorithm 1 Calculation of NBC
Input: Neural network $f$, input $\boldsymbol{x}$, neighbor input $\boldsymbol{x}^{\prime}$
Output: Neural behavior consistency between $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$
1:  $s\leftarrow 0$
2:  for the $i$-th layer $l_{i}$ in hidden layers do
3:     $\boldsymbol{v}\leftarrow[f^{(i)}(\boldsymbol{x})_{1},f^{(i)}(\boldsymbol{x})_{2},\dots,f^{(i)}(\boldsymbol{x})_{m[i]}]^{\text{T}}$
4:     $\boldsymbol{v}^{\prime}\leftarrow[f^{(i)}(\boldsymbol{x}^{\prime})_{1},f^{(i)}(\boldsymbol{x}^{\prime})_{2},\dots,f^{(i)}(\boldsymbol{x}^{\prime})_{m[i]}]^{\text{T}}$
5:     $nbc\leftarrow\frac{\boldsymbol{v}\cdot\boldsymbol{v}^{\prime}}{|\boldsymbol{v}|\cdot|\boldsymbol{v}^{\prime}|}$
6:     $s\leftarrow s+\frac{nbc}{\gamma[i]}$
7:  end for
8:  $s\leftarrow s-\textsc{KL}(f(\boldsymbol{x})||f(\boldsymbol{x}^{\prime}))$
9:  return $s$
During the calculation of NBC, the NBC value $s$ is initially set to zero, then iterating over the hidden layers of $f$, the algorithm assesses the consistency for each neuron between the original and adversarial images, incrementally updating NBC. The final NBC is calculated as the sum of the scaled neuron consistencies across all layers.
Due to the characteristics of gradient backpropagation, the layer close to the output layer has a greater impact on the network’s behavior. Therefore, we use the KL divergence as a consistency metric for the output layer as a regularization term to ensure that the network’s output remains consistent across different inputs, which can be calculated as:
$\textsc{KL}(f(\boldsymbol{x})||f(\boldsymbol{x}^{\prime}))=\sum_{i}f(\boldsymbol{x})_{i}\log\frac{f(\boldsymbol{x})_{i}}{f(\boldsymbol{x}^{\prime})_{i}}.$ (6)
For the hidden layers, to prevent gradient explosion or vanishing, we use cosine similarity as a continuous metric to approximate the consistency of behavior, in which the output is in the range of [0,1], making it more suitable for training neural networks with high-dimensional intermediate layers. To balance the impact of layers with different numbers of neurons, we scale the NBC value by driving $\gamma[i]$ at line 6 of Algorithm 1. Our intuition is to prioritize layers with smaller dimensionsas their consistent behavior may may constrain subsequent layers through propagation. Moreover, input and output layers often have fewer neurons, applying constraints to them can directly affect the forward and backward propagation process and accelerate the convergence of target loss. In contrast, over-constraining middle layers, which have more neurons and extract complex features, could limit the model’s expressive power. Roughly speaking, using a looser penalty on layers with more neurons may slightly decrease accuracy but increase the proportion of stable neurons and improve verification speed. In our experiments, we apply factor $\gamma[i]=2^{r[i]}$ to balance the impact of layers with different numbers of neurons, where $r[i]$ means the number of neurons in the $i$-th hidden layer is the $r[i]$-th smallest number in all the numbers of neurons in hidden layers.
Algorithm 2 Calculation of $l^{\textsc{NBC}}$
Input: Neural network $f$, input $\boldsymbol{x}$, label $\boldsymbol{y}$,perturbation $\varepsilon$, number of perturbation steps $k$, step size $\alpha$, NBC regularization hyperparameter $\beta$
Output: Final NBC loss $l^{\textsc{NBC}}$
1:  Generate a random starting point $\boldsymbol{x}^{\prime}\in\mathcal{S}(\boldsymbol{x},\varepsilon)$
2:  for $i$ from $1$ to $k$ do
3:     $\Delta\boldsymbol{x}\leftarrow\frac{\partial\textsc{NBC}(f,\boldsymbol{x},\boldsymbol{x}^{\prime})}{\partial\boldsymbol{x}^{\prime}}$
4:     $\boldsymbol{x}^{\prime}\leftarrow\boldsymbol{x}^{\prime}-\alpha\Delta\boldsymbol{x}$
5:     $\boldsymbol{x}^{\prime}\leftarrow\text{clip}(\boldsymbol{x}^{\prime},\boldsymbol{x},\varepsilon)$
6:  end for
7:  $l^{\textsc{NBC}}\leftarrow\text{CE}(f(\boldsymbol{x}),\boldsymbol{y})-\beta\cdot\textsc{NBC}(f,\boldsymbol{x},\boldsymbol{x}^{\prime})$
8:  return $l^{\textsc{NBC}}$
Algorithm 2 outlines the NBC loss calculation process. This algorithm uses the idea of adversarial training to find an adversarial input $\boldsymbol{x}^{\prime}$ that minimizes the NBC loss between the input $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$. The algorithm starts by randomly selecting an adversarial input $\boldsymbol{x}^{\prime}$ within the $\varepsilon$-neighborhood of the original input $\boldsymbol{x}$. Subsequently, the adversarial input undergoes iterative perturbation over $k$ steps, with each perturbation step adjusting the adversarial input by gradients of the NBC for the adversarial input, scaled by the step size $\alpha$. The adversarial input is clipped after each perturbation step to ensure that it remains within the perturbation range allowed from the original input. The final NBC loss is calculated as the summation of the cross-entropy loss and the negation of scaled NBC (Algorithm 1) between the original input and the adversarial input. Detailed disgussion of the hyperparameters $\gamma$ in Algorithm 1 and $\beta$ in Algorithm 2 are provided in supplementary material.
We train the network to maximize neural response uniformity across varied inputs within the neighborhood, thereby reducing the number of unstable neurons and tightening neuronal bounds. Consequently, the network trained with NBC loss is more friendly to formal verification methods, as it exhibits fewer unstable neurons and more precise bounds.
Evaluation
In this section, we evaluate our methods to answer the following research questions:
RQ1: Can networks trained with our method maintain their verification-friendly properties across various network architectures and radii?
RQ2: Can our method be effectively integrated with existing training methods?
RQ3: How does our method compare with existing methods under close accuracy ranges?
Figure 2: Overview of last epoch results. For each metric at each perturbation radius, the best-performing value is defined as 100, with other values scaled proportionally. We average all the scaled values of each model at each perturbation radius in each dataset to obtain the final score. For each metric, the larger value indicates better performance.
Model M1 M2 M3
Method T M R O T M R O T M R O
Test Acc. 99.0 98.2 98.9 98.8 99.1 98.3 99.2 98.9 99.1 99.3 99.3 99.0
$\varepsilon$=0.1 UNSAT96.6 95.0 96.4 97.2 96.8 93.0 96.9 97.0 94.5 87.8 86.0 96.6
Stable72.0 84.7 68.0 84.7 69.4 81.5 65.9 86.6 78.4 69.3 48.4 92.1
Time (s) 3.6 4.7 4.0 3.0 3.7 5.2 3.6 3.2 9.2 19.0 20.6 4.0
[…]
$\text{PGD}_{100}$ 97.6 95.9 97.7 97.7 97.8 94.3 97.9 97.9 98.0 97.6 98.2 97.9
$\varepsilon$=0.2 UNSAT89.5 86.7 79.0 92.2 85.6 71.3 71.8 91.8 2.3 0.2 1.5 46.3
Stable34.8 65.7 25.2 64.5 27.7 66.1 20.5 68.0 17.2 10.0 4.3 68.7
Time (s) 13.7 12.8 28.0 7.3 20.8 30.2 39.1 8.7 111.8 113.0 113.1 68.4
[…]
$\text{PGD}_{100}$ 95.7 91.8 95.0 95.6 96.0 88.8 95.8 96.4 96.1 95.5 96.0 96.0
$\varepsilon$=0.3 UNSAT29.4 35.5 7.0 60.1 13.3 8.2 3.9 47.4 0.0 0.0 0.0 0.0
Stable6.2 40.7 2.4 47.8 2.9 37.2 1.4 53.9 0.1 0.0 0.0 52.5
Time (s) 81.1 64.3 94.3 47.8 95.4 84.5 99.8 63.5 108.4 106.6 108.3 107.7
[…]
$\text{PGD}_{100}$ 93.1 83.3 89.3 91.3 92.8 80.0 91.8 91.4 92.2 91.3 91.6 91.9
Table 1: Networks trained with $\varepsilon=0.3$ on MNIST datasets. The best results are highlighted in bold. Verified under $\varepsilon$=0.1, 0.2, 0.3. T:TRADES, M:Madry, R:ReLU Stable, O:Ours.
Model M1 M2 M3
Method T* M* R* T* M* R* T* M* R*
Test Acc. -0.1 -3.9 -3.5 -0.1 -5.7 -2.8 -0.1 -0.4 -0.5
$\varepsilon$=0.1 UNSAT+0.6 -7.6 -4.3 -0.2 -11.2 -3.8 +2.0 +9.0 +10.6
Stable+12.7 +12.8 +24.4 +16.4 +15.8 +25.6 +13.7 +25.3 +44.4
Time (s) -0.7 -1.1 -0.7 +1.1 +0.9 -0.3 -3.4 -14.8 -16.3
[…]
$\text{PGD}_{100}$ +0.1 -6.1 -4.3 -4.0 -7.9 -3.5 -0.2 +0.3 -0.4
$\varepsilon$=0.2 UNSAT+2.7 -10.0 +9.4 +4.9 -11.8 +18.0 +43.0 +83.2 +87.0
Stable+29.7 +28.1 +57.9 +37.9 +26.2 +60.1 +51.0 +70.6 +74.6
Time (s) -6.4 -6.8 -23.9 -7.2 -12.3 -34.7 -41.9 -88.0 -97.3
[…]
$\text{PGD}_{100}$0.2 -9.9 -4.2 -2.2 -12.8 -4.1 -0.2 +0.4 -0.3
$\varepsilon$=0.3 UNSAT+30.7 +16.8 +72.4 +30.4 -0.6 +75.1 0.0 +4.4 +38.9
Stable+41.6 +46.2 +69.1 +47.7 +39.5 +64.9 +51.7 +57.4 +53.8
Time (s) -33.3 -44.1 -86.2 -27.2 -32.6 -90.6 -0.5 -3.9 -40.8
[…]
$\text{PGD}_{100}$1.5 -15.2 -3.0 +1.1 -21.6 -4.3 +0.2 +1.0 -0.4
Table 2: Networks trained with each method combined with our method on the MNIST dataset with $\varepsilon=0.3$. Improved results are highlighted in bold. T*: TRADES+Ours, M*: Madry+Ours, R*: ReLU Stable+Ours.
Experimental Setup
Experiments are conducted on a server with 128 Intel Xeon Platinum 8336C CPUs, 128GB memory, and four NVIDIA GeForce RTX 4090 GPUs, running Debian GNU/Linux 10 (Buster). We use Python 3.11.7 and PyTorch 2.1.2 for implementation. Other settings are as follows.
Dataset. Networks are trained on three widely used datasets, MNIST (LeCun et al. 1998), Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), and CIFAR-10 (Krizhevsky 2009).
Network Architecture. We select neural networks of varying sizes from VNN-COMP (Bak, Liu, and Johnson 2021; Müller et al. 2023) to assess the effectiveness of the methods used. For the MNIST and Fashion-MNIST datasets, we chose the M1 (cnn_4_layer), M2 (relu_stable), and M3 (conv_big) models, with approximately 0.16M, 0.17M, and 1.9M parameters, respectively. For the CIFAR-10 dataset, we select the C1 (marabou_medium), C2 (marabou_large), and C3 (conv_big) models, containing about 0.17M, 0.34M, and 2.4M parameters, respectively. Network architectures are detailed in supplementary material.
Baselines. We choose Relu Stable (Xiao et al. 2019) as a baseline, which shares the most similarities with our approach. TRADES (Zhang et al. 2019b) and Madry (Madry et al. 2018), two commonly used adversarial training methods, are selected to show that directly using our method can at least achieve robustness comparable to classical robust training methods and combining our method with existing methods can improve verification performance.
Training. We set the batch size to 128, the learning rate to 0.001, and employing the Adam optimizer. For RQ1, networks are trained under default settings for 400 epochs. For RQ2, each network underwent 200 epochs of training with the original method followed by another 200 epochs using a combination of the original and our methods, or in the reverse order. For RQ3, we use CE loss to train a base model and further fine-tune using RS, Madry, Trades, and our method separately to maintain each method’s accuracy within a specified range.
Verification. We use $\alpha,\beta$-CROWN , a state-of-the-art verification tool that performs overall best in VNN-COMP competitions, to verify the properties of the trained networks. For each dataset, we select $k$ images from each of the 10 categories in the test set. For each image $\boldsymbol{x}$, and ground truth label $y$, we verified the property that the network’s output label remain $y$ for input $x$ under each perturbation $\varepsilon$. We set $k=100$ and timeout at 120 seconds for MNIST and Fashion-MNIST, and $k=20$ and timeout at 180 seconds for CIFAR-10.
Metrics. The metrics used in our evaluation are as follows:
  • UNSAT%: The percentage of properties that are verified as holds (UNSAT), providing insight into the network’s overall verification effectiveness and verified robustness.
  • Stable%: The average percentage of stable neurons, calculated as: $\frac{\sum_{1}^{N}s_{i}}{N}$, where $s_{i}$ is stable neuron ratio of the $i$-th property, and $N$ is the total number of properties.
  • Time: The average time required to verify a property.
[…]
  • $\text{PGD}_{100}$: The network’s accuracy under a PGD attack for 100 steps with a perturbation $\varepsilon$.
Results of the Last Epoch
[…]
We observe that our method outperforms others in terms of verification time, stable neuron ratio, and verified ratio (UNSAT%). While our method slightly lags in accuracy and PGD accuracy, it generally maintains comparable accuracy to other methods. This result is a comprehensive evaluation across various perturbation radii and models, indicating that networks of various architectures trained using our method are more verification-friendly at different perturbation radii than those trained using other methods.
Table 1 shows the detailed results of networks trained on MNIST. Detailed results of Fashion-MNIST and CIFAR-10 are provided in the supplementary material. As shown in Table 1, traditional adversarial training methods enhance network robustness, yet verifiability declines as the perturbation radius increases. This indicates the need for new training techniques that support effective verification. Regarding UNSAT%, our method consistently outperforms others, especially at higher perturbation radii. For instance, when $\varepsilon=0.2$, our verified ratio reached 46.3%, which is more than 20 times that of the next best-performing method and over 231 times that of the least effective method on the M3 model. Unlike other methods, where UNSATsignificantly drops as the radius increases, our method maintains a high verified ratio across all tested radii.
As for stable neurons, our method excels across nearly all models and perturbation settings. Notably, in the M3 model at $\varepsilon=0.3$, while competing methods exhibited virtually no stable neurons, our method preserved a stable neuron ratio above 50%. A higher proportion of stable neurons means if the verification time is increased, our method is more likely to be verified (as the search space is significantly reduced).
Regarding average verification time, including UNSAT and Timeout issues, our method generally requires less time across all verification tasks. An exception occurs in the M3 model at a 0.3 perturbation radius, where verification failed to confirm all the properties of our model. Conversely, the model trained using the Madry has many properties quickly verified as violated, thus shortening the times. Furthermore, under a 100-step PGD attack, our method achieved comparable accuracy to these adversarial training approaches. Such results also hold for Fashion-MNIST and CIFAR-10 datasets (see supplementary material).
{custombox}
Answer RQ1: Our method maintains high stable neuron ratios, UNSAT%, and robustness across various perturbation radii. Overall, networks trained with our method maintain their verification-friendly properties across various network architectures and radii.
Combination with Existing Methods
Table 2 shows the results of networks trained with our method combined with other methods on MNIST dataset. All examples show that our method combined with existing methods improves the ratio of stable neurons. In most cases, our method combined with existing methods increases the ratio of verified properties (UNSAT%) and reduces the time required for verification greatly.
It is worth noting that when the baseline method is combined with our method, the performance is particularly outstanding in larger network structures. For example, in the M3 model. Our method combined with Madry and RS methods respectively increased the verified ratio by 83.2% and 87.0% at $\varepsilon=0.2$ and also decreased the time by 94.4 seconds and 102.5 seconds. This means that our method makes it possible to verify larger network structures. The same results are also reflected in the Fashion-MNIST and CIFAR-10 datasets provided in the supplementary material.
Combining our method with existing methods sacrifices some accuracy and adversarial accuracy, but improves verification and neural stability. In particular, the improvement is more pronounced in larger models. This means that our method makes it possible to verify larger network structures using existing verification tools. The same results are also reflected in the Fashion-MNIST and CIFAR datasets provided in the supplementary material.
{custombox}
Answer RQ2: Our method combined with other methods demonstrates to improve the verification-friendly properties of the network, however, a trade-off between accuracy and verifiability is required.
Comparison under Close Accuracy
Figure 3: Networks trained with $\varepsilon={2}/{255}$ on CIFAR-10 dataset and evaluated on $\varepsilon={2}/{255}$. Methods unable to achieve given accuracy range are omitted. ’Natural’ means the network are trained with only cross-entropy loss.
The accuracy of networks trained using default parameters varied significantly on the CIFAR-10 dataset. To compare the performance of networks with close accuracy, we fine-tuned the parameters to ensure that the accuracy of each method fluctuated within a certain range.
Figure 3 shows the results of networks trained on CIFAR-10. Larger perturbation radii are very challenging for current verification methods, so we choose a smaller perturbation radius to train and evaluate the networks.
In terms of UNSAT%, our method performs best in different network models and different accuracy ranges and generally requires less verification time for timeouts and UNSAT results than other methods, indicating that our networks are easier to verify. When the models reach higher accuracy, our method has a slightly longer average time, as other methods are more easily to find counterexamples using attack methods, while our method is less vulnerable to attacks.
Our accuracy under PGD is generally higher than other methods, with a few exceptions, where it is slightly lower than the best-performing method. Our method consistently maintains a higher ratio of stable neurons than other methods, especially in the C3 model, our method maintains a high ratio of stable neurons of up to 80.1% while maintaining a high verification accuracy, indicating that the upper bound of the search space is significantly reduced.
{custombox}
Answer RQ3: Experiments show that our method is more verification-friendly than other methods at higher and close accuracy levels, as our method consistently maintains a high ratio of stable neurons, and relative robustness and requires less verification time.
Related Work
Traditional adversarial methods such as TRADES (Zhang et al. 2019b) and Madry et al. (Madry et al. 2018) use adversarial training to improve the robustness of neural networks. However, these methods most impose constraints on the network output, while our method imposes neuron behavior constraints on all layers of the network. Recent works, HYDRA (Sehwag et al. 2020) integrates the pruning process with adversarial training, using a criterion that considers adversarial robustness during the pruning decision, leading to more compact yet robust models and Wu et al. (Wu, Xia, and Wang 2020) found that adversarial weight perturbation can improve the robustness of the network.
Certified training (De Palma et al. 2022; Mirman, Gehr, and Vechev 2018; Jovanovic et al. 2022; Zhang et al. 2019a; Xu et al. 2020) introduces interval bound propagation (IBP) into the training process to improve the robustness of the network but suffers from the long training time.
The ReLU Stable method (Xiao et al. 2019) ensures that the upper and lower bounds of the neurons have the same sign, pushing the upper and lower bounds away from zero in the same direction. This method can also incorporate ternary loss to train a SAT-friendly network (Narodytska et al. 2019). In contrast, our method emphasizes the stability of neurons before and after perturbation, focusing on the similarity and consistency of neuron behavior. Furthermore, RS loss requires additional calculations for each neuron’s bounds, whereas our method does not require extra information, making it easier to implement. Linearity grafting (Chen et al. 2022c) replaces unstable neurons with linear neurons to improve the robustness of the network. However, this method will modify network architectures, while our method keeps the architecture unchanged. MILP-based method (Baninajjar, Rezine, and Aminifar 2023) uses MILP to post-process the network to make it sparse to improve the verification efficiency, which may be time-consuming. Pruning-based methods (Xiao et al. 2019; Xu et al. 2024) heuristically prune the neurons that are inactive or unstable and have little impact on the network’s performance. Bias shaping method (Xu et al. 2024) changing the bias of the unstable neurons during training to stabilize them. These methods are either post-processing methods or training tricks that can be incorporated with other methods to further improve the verifiability of networks.
Conclusion
In this work, we propose a novel training method to train a verification-friendly neural network by preserving the neuron behavior consistency of the network. Our experiments on the MNIST, Fashion-MNIST, and CIFAR-10 datasets demonstrate that our method consistently enhances the network’s verification-friendliness, as evidenced by a stable neuron ratio, comparable robustness, and faster verification speed across different perturbation radii. Additionally, existing methods, when combined with ours, show improved verification efficiency. Our method also speeds up the verification process while maintaining the accuracy of the model, which is not commonly achieved by existing methods.
Acknowledgements
This work is partly supported by CAS Project for Young Scientists in Basic Research, Grant No.YSBR-040, ISCAS New Cultivation Project ISCAS-PYFX-202201, ISCAS Basic Research ISCAS-JCZD-202302 and the Ministry of Education, Singapore under its Academic Research Fund Tier 3 (Award ID: MOET32020-0004).